Use the discriminant to determine how many roots the graph of the following equation will have.
x^2-8x+16

Answer by stanbon(75887)

Use the discriminant to determine how many roots the graph of the following equation will have.
x^2-8x+16
a = 1; b = -8; c = 16
----
D = b^2-4ac = 64 -4*1*16 = 0
so there are 2 equal Real number solutions.
---------------------
x^2-8x+16 = (x-4)^2
Roots are x=4 with multiplicity 2.
================================
Cheers,
Stan H.